OpenStudy (anonymous):

What is the equation of the line through the point (-2, 3) with m = 0? Answer y = -2 y = -2x y = 3x y = 3 1 points Question 35 Write an equation for the line in slope-intercept form. Given that a line passes through the points (3, 3) and (2, 5). Answer y = -2x – 3 y = 2x + 3 y = -2x + 9 y = 2x – 9

OpenStudy (texaschic101):

if the slope of the line is 0, then it is a horizontal line....therefore y = 3 is your equation. find the slope : slope(m) = (y2 - y1) / (x2 - x1) (3,3) -- x1 = 3, y1 = 3 (2,5) -- x2 = 2, y2 = 5) now plug in your values and find the slope

OpenStudy (anonymous):

y = 3

OpenStudy (anonymous):

y = 2x – 9

OpenStudy (anonymous):

thank you!![: